Mike is 5 years less than three times as old as his sister. Next year he will be twice as old as she will be. How old is each?
Answer by TimothyLamb(4379)

---
x = Mike
y = sister
---
x = 3y - 5
x + 1 = 2(y + 1)
---
put the system of linear equations into standard form:
---
x - 3y = -5
x - 2y = 1
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = Mike = 13
y = sister = 6
---
system matrix
1 -3
1 -2
---
inverse of system matrix
-2 3
-1 1
---
determinant of system matrix = 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
)